OpenStudy (anonymous):

need help 3 questions will medal A certain drug is made from only two ingredients: compound A and compound B. There are 5 milliliters of compound A used for every 3 milliliters of compound B. If a chemist uses 246 milliliters of compound B, how many milliliters of the drug will be made?

OpenStudy (hellokitty17):

3:2 ration of A to B so out of the 5 total parts (3+2=5), 3/5 are chem a 2/5 are chem b note this adds up to 5/5 or 100% of the final product so 3/5*460 = parts of soltuion that is chem A

OpenStudy (hellokitty17):

here sorry this is the right one use the unitary method the ratio of A to B is 5:6 Compound B find 528/6 = Answer then its 5 x answer to find compound A *** to the total amount = Compound A + Compound B *** above + 528

OpenStudy (anonymous):

656 was the answe they wanted

OpenStudy (hellokitty17):

is that even an answer choice?

OpenStudy (anonymous):

there was no choices i figured they way u showed and it gave me that as the right answer